A 9-year bond has a yield of 4% and a duration of 8.286 years. If the market yield changes by 40 basis points, what is the percentage change in the bond’s price? (Do not round intermediate calculations. Negative value should be indicated by a minus sign. Round your answer to 2 decimal places.)
Answer #1

We know AB -D* Ay B1+y
percentage change in the bond’s price = -8.286 + 0.004 * * 100 = 1.04 -3.19%。

Please do rate me and mention doubts, if any, in the comments section.